A fire hose ejects a stream of water at an angle of 30.1 ° above the horizontal. The water leaves the nozzle with a speed of 25.8 m/s. Assuming that the water behaves like a projectile, how far from a building should the fire hose be located to hit the highest possible fire?
Answer #1

consider the motion along the Y-direction from initial point of launch to point A

vfy = final velocity at the highest point = 0 m/s

voy = initial velocity = vo Sin\theta = 25.8 Sin30.1 = 12.94 m/s

a = acceleration = - 9.8 m/s2

t = time for the stream to reach the maximum height = ?

using the equation

vfy = voy + at

0 = 12.94 + (- 9.8) t

t = 1.32 sec

consider the motion along the X-direction :

X = horizontal distance travelled by stream before hitting the maximum height = ?

vox = initial velocity in X-directon = vo Cos\theta  = 25.8 Cos30.1 = 22.32 m/s

t = time for the stream to reach the maximum height = 1.32 sec

ax = acceleration along the horizontal direction = 0 m/s2

Using the equation

X = vox t + (0.5) a t2

inserting the values

X = (22.32) (1.32) + (0.5) (0) (1.32)2

X = 29.5 m